OpenStudy (anonymous):

Solve the equation. c/4 - 5=4

rishavraj (rishavraj):

is it ?? \[\frac{ c }{ 4 }~-~5 = 4\]

OpenStudy (anonymous):

Yes @rishavraj

rishavraj (rishavraj):

so u need to isolate the variable.... so first of all add 5 on both sides

rishavraj (rishavraj):

- 5 + 5 = 0

rishavraj (rishavraj):

so it would be \[\frac{ c }{ 4 }~=~9\]

rishavraj (rishavraj):

@rainbowswirl there???

OpenStudy (anonymous):

Here, but confused @rishavraj

rishavraj (rishavraj):

hmmmm whts kinda confusing u ...lemme know :))

OpenStudy (anonymous):

Why did you do -5 + 5? @rishavraj

rishavraj (rishavraj):

i meant to say tht when u add 5 on both sides it would be \[\frac{ c }{ 4 } - 5 + 5 = 4 + 5\] which means \[\frac{ c }{ 4 } = 9\] u got tht now>??

OpenStudy (anonymous):

ok, thank you, is that all? @rishavraj

rishavraj (rishavraj):

now u need to get rid of 4 from c.....wht do u think u need to do to get tht

OpenStudy (anonymous):

I dont know :| @rishavraj

rishavraj (rishavraj):

hmmmmm jst asking hav u been taught this in school ????@rainbowswirl

rishavraj (rishavraj):

now jst multiply by 4 on both sides.... u see \[a \times \frac{ 1 }{ a } = 1\]

rishavraj (rishavraj):

@rainbowswirl ???

OpenStudy (anonymous):

Ok @rishavraj

rishavraj (rishavraj):

on multiplying c = 36 :))

OpenStudy (anonymous):

Thank you! @rishavraj

OpenStudy (anonymous):

@MissSmartiez is there anyway you can explain this all in one??

OpenStudy (misssmartiez):

Hm, I am doing a test, and besides, I just am keen in the basics of math, not my best subject, AT ALL. I am good in English and science rather than math. Though, he did explain it quite well. http://www.quickmath.com/webMathematica3/quickmath/equations/solve/basic.jsp#c=solve_stepssolveequation&v1=+c%2F4+-+5%3D4 Click the solution's answer; it will show you.

OpenStudy (anonymous):

Thank you! @MissSmartiez

OpenStudy (misssmartiez):

welcome :)